During 2016, Cheyenne Corporation spent $149,760 in research and development costs. As a result, a new product called the New Age Piano was patented. The patent was obtained on October 1, 2016, and had a legal life of 20 years and a useful life of 10 years. Legal costs of $23,760 related to the patent were incurred as of October 1, 2016.

Prepare all journal entries required in 2016 and 2017 as a result of the transactions above. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Date

Account Titles and Explanation

Debit

Credit

2016

enter an account title to record research and development expenses

enter a debit amount

enter a credit amount

enter an account title to record research and development expenses

enter a debit amount

enter a credit amount

(To record research and development expenses)

2016

enter an account title to record legal expenses

enter a debit amount

enter a credit amount

enter an account title to record legal expenses

enter a debit amount

enter a credit amount

(To record legal expenses)

2016

enter an account title to record amortization expense

enter a debit amount

enter a credit amount

enter an account title to record amortization expense

enter a debit amount

enter a credit amount

(To record amortization expense)

2017

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

On June 1, 2018, Cheyenne spent $9,000 to successfully prosecute a patent infringement suit. As a result, the estimate of useful life was extended to 12 years from June 1, 2018. Prepare all journal entries required in 2018 and 2019. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Date

Account Titles and Explanation

Debit

Credit

2018

enter an account title to record patents

enter a debit amount

enter a credit amount

enter an account title to record patents

enter a debit amount

enter a credit amount

(To record patents)

2018

enter an account title to record amortization expense

enter a debit amount

enter a credit amount

enter an account title to record amortization expense

enter a debit amount

enter a credit amount

(To record amortization expense)

2019

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

In 2020, Cheyenne determined that a competitor’s product would make the New Age Piano obsolete and the patent worthless by December 31, 2021. Prepare all journal entries required in 2020 and 2021. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amoun
Answer #1

Solution

Cheyenne Corporation

Journal Entries

  1. For transactions in 2016 and 2017:

Date

Account Titles and Explanation

Ref. No.

Debit

Credit

2016

Research and Development Expense

$149,760

Cash

$149,760

(To record research and development expense)

2016

Patent

$23,760

Cash

$23,760

(To record cost of patent/legal expenses)

2016

Amortization Expense – Patent

$594

Patent

$594

(To record amortization expense; 23,760/10 x 3/12 = $1,026)

2017

Amortization Expense – Patent

$2,376

Patent

$2,376

(To record amortization expense; 23,760/10 = $2,376)

  1. Journal entries for transactions in 2018 and 2019:

Date

Account Titles and Explanation

Ref. No.

Debit

Credit

2018

Patents

$9,000

Cash

$9,000

(To record patents)

2018

Amortization Expense – Patents

$2,390

Patents

$2,390

(To record amortization expense)

2019

Amortization Expense – Patents

$2,400

Patents

$2,400

(To record patents)

Computations:

Patent amortization –

For 2018 Jan 1 – June 1 (5 months) = 23,760/10 x 5/12 = $990

Book value up to June 1 = 23,760 – (594 + 2,376 + 990) = $19,800

Add: infringement cost = $9,000

Book value, patents June 1 = $28,800

Patent amortization June 1- Dec 31 (7 months) = 28,800/12 x 7/12 = $1,400

Total amortization expense, 2018 = 990 +1,400 = $2,390

Book value Jan 1, 2019 = 28,800 – 1,400 = 27,400

  • 2019 patent amortization expense = 28,800/12 = $2,400
  1. Journal entries for 2020 and 2021:

Date

Account Titles and Explanation

Ref. No.

Debit

Credit

2020

Amortization Expense – Patents

$12,500

Patents

$12,500

(To record amortization expense)

2021

Amortization Expense – Patents

$12,500

Patents

$12,500

(To record amortization expense)

Computations:

For 2020 book value -

28,800 – 1,400 – 2,400 = $25,000

Since the patent is determined to become obsolete by December 31, 2021, the entire book value would be written down in the two years, 2020 and 2021, respectively.